# Large Language Models as Visual Cross-Domain Learners

### 06 Jan 2024 v1

@@ Abstract
Recent advances achieved by deep learning models rely on the independent and identically distributed assumption, hindering their applications in real-world scenarios with domain shifts. To address the above issues, cross-domain learning aims at extracting domain-invariant knowledge to reduce the domain shift between training and testing data. However, in visual cross-domain learning, traditional methods concentrate solely on the image modality, neglecting the use of the text modality to alleviate the domain shift. In this work, we propose Large Language models as Visual cross-dOmain learners (LLaVO). LLaVO uses vision-language models to convert images into detailed textual descriptions. A large language model is then finetuned on textual descriptions of the source/target domain generated by a designed instruction template. Extensive experimental results on various cross-domain tasks under the domain generalization and unsupervised domain adaptation settings have demonstrated the effectiveness of the proposed method.
@@ 1 Introduction
Typical cross-domain learning scenarios include domain generalization (DG) [5, 75, 64] and unsupervised domain adaptation (UDA) [97, 85, 36, 95]. Specifically, DG trains a model from multiple source domains and evaluates it on an unseen target domain, while UDA assumes some unlabeled data from the target domain are available for training. Various DG [6, 40, 104] and UDA [85, 97, 45] methods have been proposed to alleviate the domain shift by learning domain-invariant features in the image modality. For example, under the DG setting, MIRO [6] proposes to learn features that are similar to the oracle representation. For the UDA setting, CDTrans [85] proposes a weight-sharing triple-branch transformer framework to enhance feature alignment and improve robustness to noisy pseudo-labels. However, the above methods only utilize image modality for cross-domain learning.
@@
Instead of just using the image modality, recent efforts on large-scale vision-language models (VLMs) [55, 103, 21] have shown significant improvements in image classification by learning from massive paired image-text samples. For example, CLIP [55], which consists of an image encoder and a text encoder, is trained on 400 million image-text pairs by contrastive learning [8]. However, as shown in Figure 0(a), learning domain-invariant features in CLIP is challenging due to the large domain shift, leading to sub-optimal cross-domain performance. Large language models (LLMs) [70, 4] have impressive zero-shot capability for text-based tasks. Nevertheless, it is essential to acknowledge that the success of LLMs does not readily extend to pure vision and vision-language tasks, primarily due to the inherent disparities in modalities and task structures. To the best of our knowledge, there is no work that integrates LLMs with visual cross-domain learning.
@@
In this paper, we propose LLaVO, using Large Language models as Visual cross-dOmain learners, to integrate the powerful LLMs into visual cross-domain learning. Specifically, we first use various VLMs (including CLIP [55] and BLIP [39]) to convert images into detailed textual descriptions (i.e., tags, attributes, and captions). As shown in Example 4.4.3 placed in the experimental section, these descriptions contain crucial image content information. As a byproduct, based on these descriptions, we release a series of datasets in text modality for text classification tasks and put the details in Appendix D. Next, we design an instruction template for querying the LLM to obtain images’ categories based on their descriptions. As shown in Figure 0(b), the domain shift of LLM embeddings is smaller than that of CLIP. Though LLMs perform well in natural language understanding, they may be distracted by irrelevant contexts [63, 88, 12] and the issue of hallucinations [74, 96, 67], leading to undesirable outputs. Especially in cross-domain learning scenarios, there are many domain-specific, but classification-irrelevant pieces of information [87]. To deal with these challenges, we finetune the LLM using training data comprised of question-answer pairs, where the proposed template creates questions, and the answers correspond to the class tokens of the respective images. In the UDA scenario, where the unlabeled target domain data are available, we use predictions of target domain data as pseudo-labels to finetune the LLMs for boosting performance. Moreover, LLaVO can use any off-the-shelf large-scale VLMs and LLMs. Extensive experiments are conducted on benchmark datasets under the DG and UDA settings to demonstrate that the proposed LLaVO method achieves state-of-the-art performance. Especially, compared with existing VLM-based methods, LLaVO performs better, showing the benefits of integrating LLMs into cross-domain learning.
@@
  • We are the first to integrate LLMs with visual cross-domain learning. The proposed LLaVO uses VLMs as a bridge between image and text modalities and designs a question instruction template for the LLM to predict the category of an image based on its description.
  • The proposed LLaVO can alleviate the domain shift by finetuning the LLM on the designed question instruction template to improve the instruction following ability. In UDA, we finetune the LLM with pseudo-labels to further improve the cross-domain performance.
  • Extensive experiments on diverse DG and UDA tasks demonstrate that LLaVO consistently achieves state-of-the-art performance, surpassing existing VLM-based methods.
  • To facilitate further research, we release a series of datasets in the text modality, including PACS, VLCS, OfficeHome, and DomainNet.
@@ 2 Related Work
Figure 2: An illustration of the proposed LLaVO framework for both UDA and DG. For DG, the source images are converted into the text modality through the VLMs. Then, the extracted descriptions are used to finetune an LLM. For UDA, we use the target domain samples with pseudo-labels, along with the source domain samples with ground truth labels to finetune the LLM.
@@
However, most existing DG and UDA methods focus solely on the image modality, leading to sub-optimal cross-domain effects as shown in our experiments. On the contrary, our LLaVO leverages both the image and text modalities.
@@
On the other hand, LLMs show impressive generalization capability in various natural language processing tasks, e.g., topic classification [31, 4], mathematical reasoning [80, 99, 89, 30], and question answering [4, 79, 98]. Moreover, instruction-tuning methods [79, 10, 52] further improve the zero-shot and few-shot performance of LLM. Despite these advancements, LLMs cannot be directly applied to pure vision tasks and vision-language tasks due to the inherent disparities in modalities and task structure. To address this issue, some works have utilized VLMs as a bridge between the image and text modalities. For example, Menon et al. [49] leverage an LLM (i.e., GPT-3 [4]) to obtain descriptions for each category and employ CLIP to make predictions based on both the images and the extracted textual descriptions. LENS [2] designs a set of highly descriptive visual modules based on VLMs, enabling LLMs to tackle visual tasks. Different from those works that assume that all the data are in the same domain, the proposed LLaVO focuses on visual cross-domain learning.
@@ 3 Methodology
In this section, we present the proposed LLaVO method.
@@ 3.1 Overview of The Entire Framework
Suppose there is a set of $S$ source domains $\{\mathcal{D}_{i}\}_{i=1}^{S}$, where $\mathcal{D}_{i}=\{({\bf x}_{j}^{i},y_{j}^{i})\}_{j=1}^{n_{i}}$ and $n_{i}$ denotes the number of samples in $\mathcal{D}_{i}$. Let $\mathcal{D}=\cup_{i=1}^{S}\mathcal{D}_{i}$ denote the entire source domain training data. In cross-domain learning, we aim to learn a model from source domains and then generalize the model to a target domain $\mathcal{T}$ whose data distribution differs from that of the source domains. In DG, target domain data are unavailable during the training process, while in the training process of UDA, the unlabeled target domain data are available. Here we study the single-source UDA and multi-source UDA is left for future research, making $S$ equal to 1 in UDA.
To begin with, LLaVO generates textual descriptions of images using pre-trained VLMs. After obtaining textual descriptions, we finetune an LLM with the designed question instruction template that combines these descriptions and their category labels. Finetuning enables the LLM to follow the question instruction template and focus on the domain-invariant information related to the classification. The overall framework of LLaVO is shown in Figure 2 and the entire algorithm is shown in Algorithm 1 of Appendix C. In the following sections, we introduce those steps one by one.
@@ 3.2 Generating Textual Descriptions for Images
For an image ${\bf x}$, we use VLMs to convert it into textual descriptions. Previous works [49, 86] have demonstrated that textual descriptions can serve as accurate information for classification. Following [2], we use pretrained VLMs to extract textual descriptions from three aspects: (i) Tags; (ii) Attributes; and (iii) Captions.
@@ 3.3 Question Instruction
After extracting textual descriptions from images, we can use LLMs for classification. Specifically, we design a question instruction template for querying the LLM to its category. The template $\mathbb{T}({\bf x})$ is shown in Example 3.3, where $\{\mathbb{D}({\bf x})\}$ denotes the placeholder of the textual description of ${\bf x}$ and {Category Set} is the placeholder of the candidate categories. An example of the category set is shown in Appendix B.1.
@@ 3.4 Finetuning
Zero-Shot LLM (ZS-LLM) [79] feeds $\mathbb{T}({\bf x})$ into LLMs and computes the sum of per-token log-likelihood for each category. The category with the highest log-likelihood is chosen as the predicted category. However, as shown in Figure 2(a), the performance of ZS-LLM is not satisfactory in classification. The limited performance could be due to two reasons: (1) The output of LLM is correlated to but still misaligned with the ground-truth category [100]. See Appendix A.2 for further discussion. (2) LLMs are susceptible to the influence of irrelevant context (e.g., domain-specific words) within the descriptions, which can negatively impact their classification accuracy [63].
To address the above issues and further exploit the capabilities of LLMs in classifying samples based on textual descriptions, we formulate the classification problem as a question-answer problem $(\mathbb{T}({\bf x}),y)$, where $y$, the class label of ${\bf x}$, is represented by a sequence of class tokens and denotes the true answer to the question $\mathbb{T}({\bf x})$.
@@ 3.4.2 Finetuning for UDA Tasks
To make LLaVO more suitable for UDA tasks, we propose the utilization of LLaVO in conjunction with pseudo-labeling techniques. Pseudo-labeling strategies [65, 90] have enjoyed success in UDA by leveraging the unlabeled target domain data [36, 73]. Under the UDA setting, LLaVO combines target domain samples with pseudo-labels and source domain samples with ground truth labels to finetune the LLM. Similar to Eq. (1), the pseudo-labeling-based finetuning objective of UDA is to minimize the training loss as
@@
where $\hat{y}$ denotes the pseduo-label for each target domain sample. Specifically, to obtain the pseudo-labels of target domain samples, the LLM is first finetuned using the source domain training data $\mathcal{D}$ with a loss function identical to Eq. (1). Subsequently, the fine-tuned LLM is employed to generate pseudo-labels for target domain samples.
@@ 4 Experiments
In this section, we empirically evaluate the proposed LLaVO method.
@@ 4.1 Experiments on DG Tasks
Datasets. Experiments are conducted on four DG benchmark datasets: PACS [37], VLCS [17], OfficeHome [73], and DomainNet [17]. Due to page limit, the statistics of each dataset are offered in Appendix B.1. We follow the leave-one-out cross-evaluation protocol [23]: for each task, one domain is chosen as the target domain for testing, while the remaining domains serve as source domains for training.
Baselines. The proposed method is compared with existing DG methods including (1) ERM-ResNet [72], SWAD [5], MIRO [6], and SAGM [75] that use only the image modality. (2) Zero-shot CLIP (ZS-CLIP) [55], ERM-CLIP, DPL [94], and CLIPood [64] that use both the image and text modalities. ZS-CLIP uses the pre-trained CLIP to calculate the similarity between the query image embedding and the text embedding from the category set, and then choose the highest one as the prediction. For ERM-CLIP, following [81], we add a classification head after the image encoder and finetune both components in the source domain. Implementation details are provided in Appendix B.2.
Results. Tables 1, 2, 3, and 4 show the classification accuracies in the target domain for the PACS, VLCS, OfficeHome, and DomainNet datasets, respectively. As can be seen, LLaVO consistently achieves the best average accuracy on all the datasets. Moreover, LLaVO outperforms CLIP-based methods (i.e., ZS-CLIP, ERM-CLIP, DPL, and CLIPood), showing that using LLM could help VLMs in learning domain-invariant features to boost cross-domain performance. LLaVO also has good performance for each task on all the datasets. Specifically, LLaVO achieves the best performance across all tasks on the PACS, OfficeHome, and DomainNet datasets, and excels in three of four tasks on the VLCS dataset. Those results demonstrate the effectiveness of the proposed LLaVO method.
@@
Table 1: Testing accuracies (%) of DG tasks on the PACS dataset. The best is in bold.
A C P S Avg
ERM-ResNet [72] 84.7 80.8 97.2 79.3 85.5
SWAD [5] 89.3 83.4 97.3 82.5 88.1
SAGM [75] 87.4 80.2 98.0 80.8 86.6
MIRO [6] 96.3 97.3 99.7 93.9 96.8
ZS-CLIP [55] 96.4 98.9 99.9 87.7 95.7
ERM-CLIP [72] 95.8 97.2 99.6 88.4 95.3
DPL [94] - - - - 97.3
CLIPood [64] 98.9 99.6 100.0 91.2 97.4
LLaVO 99.5 99.7 100.0 94.8 98.5
Table 2: Testing accuracies (%) of DG tasks on the VLCS dataset. The best is in bold.
C L S V Avg
ERM-ResNet [72] 97.7 64.3 73.4 74.6 77.5
SWAD [5] 99.8 63.3 75.3 79.2 79.1
SAGM [75] 99.0 65.2 75.1 80.7 80.0
MIRO [6] 98.4 64.5 80.3 83.7 81.7
ZS-CLIP [55] 99.7 61.8 70.1 73.9 76.4
ERM-CLIP [72] 97.0 68.9 82.8 85.1 83.5
DPL [94] - - - - 84.3
CLIPood [64] 99.7 68.1 79.9 90.8 84.6
LLaVO 99.9 69.9 81.3 88.1 84.8
Table 3: Testing accuracies (%) of DG tasks on the OfficeHome dataset. The best is in bold.
A C P R Avg
ERM-ResNet [72] 61.3 52.4 75.8 76.6 66.5
SWAD [5] 66.1 57.7 78.4 80.2 70.6
SAGM [75] 65.4 57.0 78.0 80.0 70.1
MIRO [6] 82.4 72.2 89.1 89.3 83.3
ZS-CLIP [55] 80.7 64.6 86.3 88.0 79.9
ERM-CLIP [72] 80.0 67.6 85.8 87.6 80.3
DPL [94] - - - - 84.2
CLIPood [64] 88.0 74.0 92.8 93.2 87.0
LLaVO 90.0 86.7 97.4 95.1 92.3
Table 4: Testing accuracies (%) of DG on the DomainNet dataset. The best is in bold.
C I P Q R S Avg
ERM-ResNet [72] 58.1 18.8 46.7 12.2 59.6 49.8 40.9
SWAD [5] 66.0 22.4 53.5 16.1 65.8 55.5 46.5
SAGM [75] 64.9 21.1 51.5 14.8 64.1 53.6 45.0
MIRO [6] 78.9 46.3 68.9 19.3 80.8 69.7 60.7
ZS-CLIP [55] 70.7 49.1 66.4 14.8 82.7 63.1 57.8
ERM-CLIP [72] 75.1 44.7 68.1 18.4 75.7 66.2 58.0
DPL [94] - - - - - - 56.7
CLIPood [64] 77.6 54.7 72.5 20.7 85.7 69.9 63.5
LLaVO 83.7 56.7 75.1 21.3 87.8 74.4 66.5

### 16 Mar 2024 v2

# Project Page: https://ll-a-vo.github.io/

@@ Abstract
Recent advances achieved by deep learning models rely on the independent and identically distributed assumption, hindering their applications in real-world scenarios with domain shifts. To tackle this issue, cross-domain learning aims at extracting domain-invariant knowledge to reduce the domain shift between training and testing data. However, in visual cross-domain learning, traditional methods concentrate solely on the image modality, disregarding the potential benefits of incorporating the text modality. In this work, we propose VLLaVO, combining Vision language models and Large Language models as Visual cross-dOmain learners. VLLaVO uses vision-language models to convert images into detailed textual descriptions. A large language model is then finetuned on textual descriptions of the source/target domain generated by a designed instruction template. Extensive experimental results under domain generalization and unsupervised domain adaptation settings demonstrate the effectiveness of the proposed method.
Keywords:
Domain Generalization Unsupervised Domain Adaptation Large Language Models
@@ 1 Introduction
Typical cross-domain learning scenarios include domain generalization (DG) [22, 49, 19] and unsupervised domain adaptation (UDA) [16, 18, 10, 1]. Specifically, DG trains a model from multiple source domains and evaluates it on an unseen target domain, while UDA assumes some unlabeled data from the target domain is available for training. Various DG [21, 103, 105] and UDA [18, 16, 104] methods have been proposed to alleviate the domain shift by learning domain-invariant features in the image modality. For example, under the DG setting, MIRO [21] proposes to learn features that are similar to the oracle representation. For the UDA setting, BCAT [107] leverages a bidirectional cross-attention mechanism to extract implicit source and target mixup feature representations. However, the above methods only utilize the image modality for cross-domain learning.
@@
Instead of just using the image modality, recent efforts on large-scale vision-language models (VLMs) [23, 100, 56] have shown significant improvements in image classification by learning from massive paired image-text samples. For example, CLIP [23], which consists of an image encoder and a text encoder, is trained on 400 million image-text pairs by contrastive learning [101]. However, as shown in Figure 0(a), learning domain-invariant features based on CLIP is challenging due to the large domain shift, leading to sub-optimal cross-domain performance. On the other hand, large language models (LLMs) [72, 27] have impressive zero-shot capability for text-based tasks. Nevertheless, it is essential to acknowledge that the success of LLMs does not readily extend to pure vision and vision-language tasks, primarily due to the inherent disparities in modalities and task structures. To the best of our knowledge, there is no work that integrates LLMs with visual cross-domain learning.
@@
In this paper, we propose VLLaVO, combining Vision language models and Large Language models as Visual cross-dOmain learners, to integrate the powerful LLMs into visual cross-domain learning. Firstly, we use various VLMs (including CLIP [23] and BLIP [30]) to convert images into detailed textual descriptions (i.e., tags, attributes, and captions). As shown in Example 4.4.3 placed in the experimental section, these descriptions contain crucial image content information. As a byproduct, based on these descriptions, we release a series of datasets in text modality for text classification tasks and the details are given in Appendix 0.B. Next, we design an instruction template for querying the LLM to obtain images’ categories based on their descriptions. However, the domain shift of generated textual descriptions still exists across domains. A straightforward method rank classification [27] employs the output probability of LLM for each category to perform classification. Empirical results indicate that this simple method performs poorly in challenging cross-domain tasks due to distractions from irrelevant contexts [70, 95, 96] and the issue of hallucinations [53, 69, 97]. Especially in cross-domain learning scenarios, there are many domain-specific but class-irrelevant contexts [99]. To deal with these challenges, we finetune the LLM using training data comprised of question-answer pairs, where the questions are created by the proposed template and the answers correspond to the class tokens of corresponding images. To better adapt VLLaVO in the UDA scenario, we utilize the available target domain data with its pseudo-labels to further finetune the LLM. Experiments conducted on benchmark datasets under the DG and UDA settings have demonstrated that the proposed VLLaVO method achieves state-of-the-art performance. Moreover, compared with existing VLM-based methods, VLLaVO performs better, showing the benefits of integrating LLM into cross-domain learning. Additional support can be gained from the smaller domain shift of LLM than that of VLM (i.e., CLIP), as shown in Figure 1.
@@
  • We are the first to utilize LLMs for visual cross-domain learning. The proposed VLLaVO uses VLMs for bridging image and text modalities and designs a question instruction template for the LLM to predict the category of an image based on its description.
  • The introduced VLLaVO alleviates domain shifts through the inherent generalization capability of LLM. Finetuning LLM on the designed question instruction template enhances its instruction-following ability, reducing the distractions of irrelevant contexts.
  • Extensive experiments on diverse DG and UDA tasks demonstrate that VLLaVO consistently achieves state-of-the-art performance, surpassing existing VLM-based methods.
  • To facilitate further research, we release a series of datasets in the text modality, including PACS, OfficeHome, and DomainNet.
@@ 2 Related Work
Figure 2: An illustration of the proposed VLLaVO framework for both UDA and DG. For DG, the source images are converted into the text modality through the VLMs. Then, the extracted descriptions are used to finetune an LLM. For UDA, we use the target domain samples with pseudo-labels and the source domain samples with ground truth labels to finetune the LLM.
@@
However, most existing DG and UDA methods focus solely on the image modality, leading to sub-optimal cross-domain effects as shown in our experiments. On the contrary, our VLLaVO further leverages the text modalities.
@@
On the other hand, LLMs show impressive generalization capability in various natural language processing tasks, e.g., topic classification [2, 27], mathematical reasoning [78, 79, 4, 3], and question answering [27, 28, 77]. Moreover, instruction-tuning methods [28, 85, 87] further improve the zero-shot and few-shot performance of LLM. Despite these advancements, LLMs cannot be directly applied to pure vision tasks and vision-language tasks due to the inherent disparities in modalities and task structure. To address this issue, some works have utilized VLMs as a bridge between the image and text modalities. For example, Menon et al. [9] leverage an LLM (i.e., GPT-3 [27]) to obtain descriptions for each category and employ CLIP to make predictions based on both the images and the extracted textual descriptions. LENS [25] designs a set of highly descriptive visual modules based on VLMs, enabling LLMs to tackle visual tasks. Different from those works that assume that all the data are in the same domain, the proposed VLLaVO focuses on visual cross-domain learning.
@@ 3.1 Overview of The Entire Framework
Suppose there is a set of $S$ source domains $\{\mathcal{D}_{i}\}_{i=1}^{S}$, where $\mathcal{D}_{i}=\{({\bf x}_{j}^{i},y_{j}^{i})\}_{j=1}^{n_{i}}$ and $n_{i}$ denotes the number of samples in $\mathcal{D}_{i}$. Let $\mathcal{D}=\cup_{i=1}^{S}\mathcal{D}_{i}$ denote the entire source domain training data. In cross-domain learning, we aim to learn a model from source domains and then generalize the model to a target domain $\mathcal{T}$ whose data distribution differs from that of the source domains. In DG, target domain data are unavailable during the training process, while in the training process of UDA, the unlabeled target domain data are available. In UDA, $S$ equals to 1.
To begin with, VLLaVO generates textual descriptions of images using pre-trained VLMs. These obtained textual descriptions still exhibit domain shifts. Thus, we finetune an LLM with the designed question instruction template that combines these descriptions and their category labels. Finetuning enables the LLM to follow the question instruction template and focus on the domain-invariant information related to the classification. The overall framework of VLLaVO is shown in Figure 2 and the entire algorithm is shown in Algorithm 1 of Appendix 0.C. In the following sections, we will detail each step individually.
@@ 3.2 Generating Textual Descriptions for Images
For an image ${\bf x}$, we use VLMs to convert it into textual descriptions. Previous works [9, 61] have demonstrated that textual descriptions can provide information for classification. Following [25], we use pretrained VLMs to extract textual descriptions from three aspects: (i) Tags; (ii) Attributes; and (iii) Captions.
@@ 3.3 Question Instruction
After extracting textual descriptions from images, we can use LLM for classification. Specifically, we design a question instruction template for querying the LLM to its category. The template $\mathbb{T}({\bf x})$ is shown in Example 3.3, where $\{\mathbb{D}({\bf x})\}$ denotes the placeholder of the textual description of ${\bf x}$ and {Category Set} is the placeholder of the candidate categories. An example of the category set is shown in Appendix 0.A.1.
@@ 3.4 Finetuning
The question instruction template with textual descriptions effectively translates cross-domain image classification tasks into text classification tasks. Zero-Shot LLM (ZS-LLM) [28] feeds $\mathbb{T}({\bf x})$ into LLM and computes the sum of per-token log-likelihood for each category to apply rank classification [27]. The category with the highest log-likelihood is chosen as the predicted category. However, as shown in Figure 2(a), ZS-LLM performs poor in cross-domain tasks because (1) As shown in Table 8, the distribution of textual description is still different across domains. LLM are susceptible to the influence of irrelevant context (e.g., domain-specific words) within the descriptions, which can negatively impact their classification accuracy [70]. (2) The output of LLM is correlated to but still misaligned with the ground-truth category [98]. See Appendix 0.E.2 for further discussion.
To address the above issues and further exploit the capabilities of LLM in classifying samples based on textual descriptions, we formulate the classification problem as a question-answer problem $(\mathbb{T}({\bf x}),y)$, where $y$, the class label of ${\bf x}$, is represented by a sequence of class tokens and denotes the true answer to the question $\mathbb{T}({\bf x})$.
@@ 3.4.2 Finetuning for UDA Tasks
To make VLLaVO more suitable for UDA tasks, we propose the utilization of VLLaVO in conjunction with pseudo-labeling techniques. Pseudo-labeling strategies [34, 35] have enjoyed success in UDA by leveraging the unlabeled target domain data [10, 36]. Under the UDA setting, VLLaVO combines target domain samples with pseudo-labels and source domain samples with ground truth labels to finetune the LLM. Similar to Eq. (1), the pseudo-labeling-based finetuning objective of UDA is to minimize the training loss as
@@
where $\hat{y}$ denotes the pseudo-label for each target domain sample. Specifically, the LLM is first finetuned using the source domain training data $\mathcal{D}$ with a loss function identical to Eq. (1). Subsequently, the finetuned LLM is employed to generate pseudo-labels for target domain samples.
@@ 4 Experiments
In this section, we empirically evaluate the proposed VLLaVO method on both DG and UDA tasks. Section 4.1 discusses the experiments on DG tasks, and Section 4.2 discusses the experiments on UDA tasks. Additionally, to delve deeper into the capabilities of VLLaVO, we perform an ablation study in Section 4.3 and provide a detailed analysis in Section 4.4.
@@ 4.1 Experiments on DG Tasks
Table 1: Testing accuracies (%) of DG tasks on the PACS, OfficeHome, and DomainNet datasets. n/a denotes not applicable. The best is in bold.
PACS OfficeHome DomainNet
A C P S Avg A C P R Avg C I P Q R S Avg
ERM-ResNet [11] 84.7 80.8 97.2 79.3 85.5 61.3 52.4 75.8 76.6 66.5 58.1 18.8 46.7 12.2 59.6 49.8 40.9
SWAD [22] 89.3 83.4 97.3 82.5 88.1 66.1 57.7 78.4 80.2 70.6 66.0 22.4 53.5 16.1 65.8 55.5 46.5
SAGM [49] 87.4 80.2 98.0 80.8 86.6 65.4 57.0 78.0 80.0 70.1 64.9 21.1 51.5 14.8 64.1 53.6 45.0
MIRO [21] 96.3 97.3 99.7 93.9 96.8 82.4 72.2 89.1 89.3 83.3 78.9 46.3 68.9 19.3 80.8 69.7 60.7
ZS-CLIP [23] 96.4 98.9 99.9 87.7 95.7 80.7 64.6 86.3 88.0 79.9 70.7 49.1 66.4 14.8 82.7 63.1 57.8
ERM-CLIP [11] 95.8 97.2 99.6 88.4 95.3 80.0 67.6 85.8 87.6 80.3 75.1 44.7 68.1 18.4 75.7 66.2 58.0
DPL [65] - - - - 97.3 - - - - 84.2 - - - - - - 56.7
CLIPood [19] 98.9 99.6 100.0 91.2 97.4 88.0 74.0 92.8 93.2 87.0 77.6 54.7 72.5 20.7 85.7 69.9 63.5
ZS-LLM [28] 87.7 94.1 97.4 93.8 93.3 59.0 53.0 70.0 67.0 62.2 n/a n/a n/a n/a n/a n/a n/a
VLLaVO 99.5 99.7 100.0 94.8 98.5 90.0 86.7 97.4 95.1 92.3 83.7 56.7 75.1 21.3 87.8 74.4 66.5
Datasets. Experiments are conducted on four DG benchmark datasets, including PACS [20], OfficeHome [36], and DomainNet [24]. Due to the page limit, the statistics of each dataset are offered in Appendix 0.A.1. We follow the leave-one-out cross-evaluation protocol [71]: for each task, one domain is chosen as the target domain for testing, while the remaining domains serve as source domains for training.
Baselines. The proposed method is compared with existing DG methods including (i) ERM-ResNet [11], SWAD [22], MIRO [21], and SAGM [49] that use only the image modality. (ii) Zero-shot CLIP (ZS-CLIP) [23], ERM-CLIP, DPL [65], and CLIPood [19] that use both the image and text modalities. ZS-CLIP uses the pre-trained CLIP to calculate the similarity between the query image embedding and the text embedding from the category set, and then choose the highest one as the prediction. For ERM-CLIP, following [90], we add a classification head after the image encoder and finetune both components in the source domain. (iii) ZS-LLM [28] classifies images based on the textual descriptions. ZS-LLM and VLLaVO use the same VLMs (i.e., CLIP-ViT-H-14 [23] and BLIP [30]) for extracting textual descriptions and LLM (i.e., LLaMA-7B [72]) for classification. Note that the inference time of ZS-LLM grows with the number of categories, making it computationally infeasible for huge datasets like DomainNet (with 345 categories). Therefore, ZS-LLM is not applicable DomainNet. Implementation details are provided in Appendix 0.A.2.
Results. Table 1 shows the classification accuracies in the target domain for the PACS, OfficeHome, and DomainNet datasets, respectively. As can be seen, VLLaVO consistently achieves the best accuracy for each task on all the datasets. VLLaVO outperforms CLIP-based methods (i.e., ZS-CLIP, ERM-CLIP, DPL, and CLIPood), showing that using LLM could help VLMs in learning domain-invariant features to boost cross-domain performance. VLLaVO also outperforms ZS-LLM by a large margin (i.e., PACS(+5.2%), OfficeHome(+30.1%)), showing that finetuning the LLM on the designed question instruction template can improve its ability in handling vision cross-domain problems. Those results demonstrate the effectiveness of the proposed VLLaVO method.